# arXiv:2511.11225v2[cs.IT] 27 Feb 2026

### v2

@@ Appendix A Radiation Power via Poynting Vector
$\bm{e}_{\mathrm{rad}}(\mathbf{r})=-j\kappa_{0}Z_{0}\frac{e^{j\kappa_{0}R}}{4\pi R}\left(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\hat{\bm{\kappa}}\hat{\bm{\kappa}}^{T}\right)$
$\times\underbrace{\int_{\mathcal{S}}\bm{j}_{\mathrm{t}}(\mathbf{s})e^{-j\hat{\bm{\kappa}}^{T}\mathbf{s}}d\mathbf{s}}_{\triangleq\mathbf{J}_{\mathrm{t}}(\theta,\phi)},$ (93)
where $\hat{\bm{\kappa}}=\kappa_{0}\mathbf{k}(\theta,\phi)$, similar to (23). Plugging (A) into (92) and using the identity $\big(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\hat{\bm{\kappa}}\hat{\bm{\kappa}}^{T}\big)^{2}=\big(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\hat{\bm{\kappa}}\hat{\bm{\kappa}}^{T}\big)$ gives
$P_{\mathrm{rad}}$ $=\frac{\kappa_{0}^{2}Z_{0}}{32\pi^{2}}\lim_{R\rightarrow\infty}\oint_{\Omega}\left\|\left(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\hat{\bm{\kappa}}\hat{\bm{\kappa}}^{T}\right)\mathbf{J}_{\mathrm{t}}(\theta,\phi)\right\|^{2}\,d\Omega$
$\hskip-5.69046pt=\frac{\kappa_{0}^{2}Z_{0}}{32\pi^{2}}\int_{\mathcal{S}}\int_{\mathcal{S}}\bm{j}_{\mathrm{t}}^{H}(\mathbf{s})\left[\oint_{\Omega}\left(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\hat{\bm{\kappa}}\hat{\bm{\kappa}}^{T}\right)e^{j\hat{\bm{\kappa}}^{T}(\mathbf{s}-\mathbf{z})}d\Omega\right]$
@@
Using the identity $\nabla\nabla e^{j\hat{\bm{\kappa}}^{T}\mathbf{s}}=-\hat{\bm{\kappa}}\hat{\bm{\kappa}}^{T}e^{j\hat{\bm{\kappa}}^{T}\mathbf{s}}$, the inner integral can be simplified into
$\oint_{\Omega}$ $\left(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\hat{\bm{\kappa}}\hat{\bm{\kappa}}^{T}\right)e^{j\hat{\bm{\kappa}}^{T}\mathbf{s}}d\Omega$
$=\left(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\nabla\nabla\right)\oint_{\Omega}e^{j\hat{\bm{\kappa}}^{T}\mathbf{s}}d\Omega$
@@ Appendix B Proof of Proposition 1
To facilitate the derivation, we representation the coordinates by $s_{x}=r\sin\psi$ and $s_{y}=r\cos\psi$, where $r=\sqrt{s_{x}^{2}+s_{y}^{2}}=\|\mathbf{s}\|$. The function $\varphi(\mathbf{s})$ can be rewritten as $\varphi(\mathbf{s})=\sin(\kappa_{0}r)/(4\pi r)$. Following the chain rule of derivative, the second-order derivative along the $y$-axis is given by
@@ Appendix C Proof of Theorem 1
$\chi(\mathbf{s})=\left(\int_{\mathcal{S}}h(\mathbf{z})\mathbf{w}(\mathbf{z})d\mathbf{z}\right)h^{*}(\mathbf{s})-\frac{\mu}{2}\int_{\mathcal{S}}c(\mathbf{s}-\mathbf{z})w(\mathbf{z})d\mathbf{z}.$ (103)
@@
$\int_{\mathcal{S}}c(\mathbf{s}-\mathbf{z})w(\mathbf{z})d\mathbf{z}=\frac{2}{\mu}\left(\int_{\mathcal{S}}h(\mathbf{z})\mathbf{w}(\mathbf{z})d\mathbf{z}\right)h^{*}(\mathbf{s}).$ (104)
Observing the above condition of the optimal solution, the optimal $w(\mathbf{s})$ must be a scaled version of $v(\mathbf{s})$ satisfying (28), i.e., $w^{\mathrm{opt}}(\mathbf{s})=\alpha v(\mathbf{s})$. To find the scaling constant $\alpha$, we substitute this solution back into the power constraint (26b), which yields
@@
Appendix D Derivation of (32)
@@
$g(\mathbf{s})=\frac{j}{8\pi^{2}}\iint_{-\infty}^{+\infty}\frac{e^{j\bm{\kappa}^{T}\mathbf{s}}}{k_{z}(\bm{\kappa})}\,d\bm{\kappa},\quad k_{z}(\bm{\kappa})=\sqrt{\kappa_{0}^{2}-\|\bm{\kappa}\|^{2}}.$ (107)
For $\|\bm{\kappa}\|\leq\kappa_{0}$, $k_{z}(\bm{\kappa})$ is real. For $\|\bm{\kappa}\|>\kappa_{0}$, $k_{z}(\bm{\kappa})=j\sqrt{\|\bm{\kappa}\|^{2}-\kappa_{0}^{2}}$ is purely imaginary. In the latter case, $j/k_{z}(\bm{\kappa})$ is purely real and hence does not contribute to $\Im\{g(\mathbf{s})\}$. Consequently, $\varphi(\mathbf{s})$ is supported only on the disk $\|\bm{\kappa}\|\leq\kappa_{0}$, which yields
@@ Appendix E Proof of Proposition 4
$=$ $\frac{1}{P_{\mathrm{t}}}\left|\iint_{-\infty}^{+\infty}\beta w(\mathbf{s})e^{-j\bm{\kappa}_{\mathrm{r}}\mathbf{s}}d\mathbf{s}\right|^{2}=\frac{|\beta|^{2}}{P_{\mathrm{t}}}\left|W(\widetilde{\bm{\kappa}}_{\mathrm{r}})\right|^{2},$
where $\widetilde{\bm{\kappa}}_{\mathrm{r}}$ is defined by setting the last entry of $\bm{\kappa}_{\mathrm{r}}$ to zeros to match the definition of the 2D Fourier transform, i.e., $\widetilde{\bm{\kappa}}_{\mathrm{r}}=[\cos\theta\sin\phi,\sin\theta\sin\phi,0]^{T}$. Next, we transform the power constraint (26b). This requires the Fourier transform of the overall coupling kernel $c(\mathbf{s})$, which is given by
@@
$\max_{W(\bm{\kappa})}\quad$ $\left|W(\widetilde{\bm{\kappa}}_{\mathrm{r}})\right|^{2}$ (115a)
@@
$\left|W(\widetilde{\bm{\kappa}}_{\mathrm{r}})\right|^{2}=\left|\frac{1}{\sqrt{C(\widetilde{\bm{\kappa}}_{\mathrm{r}})}}\iint_{-\infty}^{+\infty}\sqrt{C(\bm{\kappa})}W(\bm{\kappa})\delta(\bm{\kappa}-\widetilde{\bm{\kappa}}_{\mathrm{r}})d\bm{\kappa}\right|^{2}$
$\leq\frac{1}{C(\widetilde{\bm{\kappa}}_{\mathrm{r}})}\underbrace{\iint_{-\infty}^{+\infty}C(\bm{\kappa})\left|W(\bm{\kappa})\right|^{2}d\bm{\kappa}}_{\leq 8\pi^{2}P_{\mathrm{t}}}\underbrace{\iint_{-\infty}^{+\infty}\left|\delta(\bm{\kappa}-\widetilde{\bm{\kappa}}_{\mathrm{r}})\right|^{2}d\bm{\kappa}}_{=1}$
$\leq\frac{8\pi^{2}P_{\mathrm{t}}}{C(\widetilde{\bm{\kappa}}_{\mathrm{r}})}.$ (116)
@@
$\lim_{|\mathcal{S}|\rightarrow+\infty}G_{\mathrm{opt}}$ $=\frac{8\pi^{2}|\beta|^{2}}{Z_{s}+C_{\mathrm{rad}}(\widetilde{\bm{\kappa}}_{\mathrm{r}})}$
$=\left(\frac{\kappa_{0}}{R_{0}}\right)^{2}\frac{Z_{0}^{2}(1-\sin^{2}\theta\sin^{2}\phi)^{2}\cos\phi}{2Z_{s}\cos\phi+Z_{0}(1-\sin^{2}\theta\sin^{2}\phi)},$ (117)

### v3

@@ Appendix A Radiation Power via Poynting Vector
$\bm{e}_{\mathrm{rad}}(\mathbf{r})=-j\kappa_{0}Z_{0}\frac{e^{j\kappa_{0}R}}{4\pi R}\left(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\mathbf{k}\mathbf{k}^{T}\right)$
$\times\underbrace{\int_{\mathcal{S}}\bm{j}_{\mathrm{t}}(\mathbf{s})e^{-j\mathbf{k}^{T}\mathbf{s}}d\mathbf{s}}_{\triangleq\mathbf{J}_{\mathrm{t}}(\theta,\phi)},$ (95)
where $\mathbf{k}=\kappa_{0}\mathbf{d}(\theta,\phi)$. Plugging (A) into (94) and using the identity $\big(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\mathbf{k}\mathbf{k}^{T}\big)^{2}=\big(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\mathbf{k}\mathbf{k}^{T}\big)$ gives
$P_{\mathrm{rad}}$ $=\frac{\kappa_{0}^{2}Z_{0}}{32\pi^{2}}\lim_{R\rightarrow\infty}\oint_{\Omega}\left\|\left(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\mathbf{k}\mathbf{k}^{T}\right)\mathbf{J}_{\mathrm{t}}(\theta,\phi)\right\|^{2}\,d\Omega$
$\hskip-5.69046pt=\frac{\kappa_{0}^{2}Z_{0}}{32\pi^{2}}\int_{\mathcal{S}}\int_{\mathcal{S}}\bm{j}_{\mathrm{t}}^{H}(\mathbf{s})\left[\oint_{\Omega}\left(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\mathbf{k}\mathbf{k}^{T}\right)e^{j\mathbf{k}^{T}(\mathbf{s}-\mathbf{z})}d\Omega\right]$
@@
Using the identity $\nabla\nabla e^{j\mathbf{k}^{T}\mathbf{s}}=-\mathbf{k}\mathbf{k}^{T}e^{j\mathbf{k}^{T}\mathbf{s}}$, the inner integral can be simplified into
$\oint_{\Omega}$ $\left(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\mathbf{k}\mathbf{k}^{T}\right)e^{j\mathbf{k}^{T}\mathbf{s}}d\Omega$
$=\left(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\nabla\nabla\right)\oint_{\Omega}e^{j\mathbf{k}^{T}\mathbf{s}}d\Omega$
@@ Appendix B Proof of Proposition 1
To facilitate the derivation, we represent the coordinates by $s_{x}=r\sin\psi$ and $s_{y}=r\cos\psi$, where $r=\sqrt{s_{x}^{2}+s_{y}^{2}}=\|\mathbf{s}\|$. The function $\varphi(\mathbf{s})$ can be rewritten as $\varphi(\mathbf{s})=\sin(\kappa_{0}r)/(4\pi r)$. Following the chain rule of derivative, the second-order derivative along the $y$-axis is given by
@@ Appendix C Proof of Theorem 1
$\chi(\mathbf{s})=\left(\int_{\mathcal{S}}h(\mathbf{z})w(\mathbf{z})d\mathbf{z}\right)h^{*}(\mathbf{s})-\frac{\mu}{2}\int_{\mathcal{S}}c(\mathbf{s}-\mathbf{z})w(\mathbf{z})d\mathbf{z}.$ (105)
@@
$\int_{\mathcal{S}}c(\mathbf{s}-\mathbf{z})w(\mathbf{z})d\mathbf{z}=\frac{2}{\mu}\left(\int_{\mathcal{S}}h(\mathbf{z})w(\mathbf{z})d\mathbf{z}\right)h^{*}(\mathbf{s}).$ (106)
Observing the above condition of the optimal solution, the optimal $w(\mathbf{s})$ must be a scaled version of $v(\mathbf{s})$ satisfying (28), i.e., $w_{\mathrm{opt}}(\mathbf{s})=\alpha v(\mathbf{s})$. To find the scaling constant $\alpha$, we substitute this solution back into the power constraint (26b), which yields
@@
Appendix D Derivation of the Wavenumber-Domain
Coupling Kernel (32)
@@ Appendix D Derivation of the Wavenumber-Domain
$g(\mathbf{s})=\frac{j}{8\pi^{2}}\iint_{-\infty}^{+\infty}\frac{e^{j\bm{\kappa}^{T}\mathbf{s}}}{\kappa_{z}(\bm{\kappa})}\,d\bm{\kappa},\quad\kappa_{z}(\bm{\kappa})=\sqrt{\kappa_{0}^{2}-\|\bm{\kappa}\|^{2}}.$ (109)
For $\|\bm{\kappa}\|\leq\kappa_{0}$, $\kappa_{z}(\bm{\kappa})$ is real. For $\|\bm{\kappa}\|>\kappa_{0}$, $\kappa_{z}(\bm{\kappa})=j\sqrt{\|\bm{\kappa}\|^{2}-\kappa_{0}^{2}}$ is purely imaginary. In the latter case, $j/\kappa_{z}(\bm{\kappa})$ is purely real and hence does not contribute to $\Im\{g(\mathbf{s})\}$. Consequently, $\varphi(\mathbf{s})$ is supported only on the disk $\|\bm{\kappa}\|\leq\kappa_{0}$, which yields
@@ Appendix E Proof of Proposition 4
$=$ $\frac{1}{P_{\mathrm{t}}}\left|\iint_{-\infty}^{+\infty}\beta_{\mathrm{r}}w(\mathbf{s})e^{-j\bm{\kappa}_{\mathrm{r}}\mathbf{s}}d\mathbf{s}\right|^{2}=\frac{|\beta_{\mathrm{r}}|^{2}}{P_{\mathrm{t}}}\left|W(\bm{\kappa}_{\mathrm{r}})\right|^{2}.$
Next, we transform the power constraint (26b). This requires the Fourier transform of the overall coupling kernel $c(\mathbf{s})$, which is given by
@@
$\max_{W(\bm{\kappa})}\quad$ $\left|W(\bm{\kappa}_{\mathrm{r}})\right|^{2}$ (117a)
@@
$\left|W(\bm{\kappa}_{\mathrm{r}})\right|^{2}=\left|\frac{1}{\sqrt{C(\bm{\kappa}_{\mathrm{r}})}}\iint_{-\infty}^{+\infty}\sqrt{C(\bm{\kappa})}W(\bm{\kappa})\delta(\bm{\kappa}-\bm{\kappa}_{\mathrm{r}})d\bm{\kappa}\right|^{2}$
$\leq\frac{1}{C(\bm{\kappa}_{\mathrm{r}})}\underbrace{\iint_{-\infty}^{+\infty}C(\bm{\kappa})\left|W(\bm{\kappa})\right|^{2}d\bm{\kappa}}_{\leq 8\pi^{2}P_{\mathrm{t}}}\underbrace{\iint_{-\infty}^{+\infty}\left|\delta(\bm{\kappa}-\bm{\kappa}_{\mathrm{r}})\right|^{2}d\bm{\kappa}}_{=1}$
$\leq\frac{8\pi^{2}P_{\mathrm{t}}}{C(\bm{\kappa}_{\mathrm{r}})}.$ (118)
@@
$\lim_{|\mathcal{S}|\rightarrow+\infty}\!\!G_{\mathrm{opt}}$ $=\frac{8\pi^{2}|\beta_{\mathrm{r}}|^{2}}{Z_{s}+C_{\mathrm{rad}}(\bm{\kappa}_{\mathrm{r}})}$
$=\left(\frac{\kappa_{0}}{R_{0}}\right)^{2}\!\!\!\!\frac{Z_{0}^{2}(1-\sin^{2}\theta_{0}\sin^{2}\phi_{0})^{2}\cos\phi_{0}}{2Z_{s}\cos\phi_{0}+Z_{0}(1-\sin^{2}\theta_{0}\sin^{2}\phi_{0})},$ (119)